OpenStudy (anonymous):

The binomial x + 4 is one factor of 3x3 + 5x2 - 26x + 8. Use complete sentences to explain how you use (either) polynomial division or synthetic division to find the remaining factors. List the remaining factors. Please help I have no clue

OpenStudy (anonymous):

actually you need neither method. you can see what you need by writing \[3x^3 + 5x^2 - 26x + 8=(x+4)(ax^2+bx+c)\] and thinking about what a, b and c have to be

OpenStudy (anonymous):

I wish I could say that but we have to use either one of those

OpenStudy (anonymous):

we know \(a=3\) since that is the only way you are going to get \(3a^3\) and we know \(c=2\) because that is the only way you are going to get 8

OpenStudy (anonymous):

well if your teacher does not like you to think, then we can use synthetic division, which is easy to do but hard to write here

OpenStudy (anonymous):

list the coefficients and write 3 5 -26 8

OpenStudy (anonymous):

then underneath put -4 3 5 -26 8 -4 ______________________

OpenStudy (anonymous):

bring down the 3 3 5 -26 8 -4 ______________________ 3

OpenStudy (anonymous):

I did out and I got 0 at the end

OpenStudy (anonymous):

then 3 times -4 is -12 3 5 -26 8 -4 -12 ____________________ 3

OpenStudy (anonymous):

but it says list the remaining factors what does that mean

OpenStudy (anonymous):

ooh so you did it? what did you get?

OpenStudy (anonymous):

I got the remaining of zero so it would be what 3x^2-7x+2?

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

??

OpenStudy (anonymous):

sorry i got logged out and couldn't get back in

OpenStudy (anonymous):

they wanted you to finish factoring as \[3x^2-7x+2=(3x-1)(x-2)\]